Keyboard Mapping for Function Keys F1 - F12

Date: June 25, 1996

Modifier F1 F2 F3 F4 F5 F6 F7 F8 F9 F10 F11 F12
None  ⋀ ⟹ ≡ ≠
Shift  ⇐ = ⇒ ⇔ ⊕ ⊖ ⊗ ⊘ ♮ ∞ ≤ ≥
Alt  ← − → ↔ ↠ ↦ ⇝ ↑ ↓ ∅  ∈  ∉
AltGraph ∩ ∪ ⋂ ⋃ ⊆ □ ◇ ∘ • ∥  √ \begin{matrix}c \\ ◯ \end{matrix}
Ctrl  ⊓ ⊔ \overline{| |} ⨆ ⌈ ⌉ ⌊ ⌋ (| |)  [[  ]]

Meta